OpenStudy (anonymous):

(03.06 MC) Choose the equation below that represents the line passing through the point (2, – 5) with a slope of −3. y = −3x − 13 y = −3x + 11 y = −3x + 13 y = −3x + 1

OpenStudy (anonymous):

@AnitaS.

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

this one i have no clue what it is

OpenStudy (anitas.):

any guesses?

OpenStudy (anonymous):

none because i dont know how they got 11 or 13

OpenStudy (anitas.):

I would say it is D the last one because it shows when you do the work like this. y=mx+b y=(-3) · x+b y=(-3) · (2)+b (-5)=(-3) · (2)+b (-3) · (2)+b=-5 (-3)(2)+b=-5 (-6)+b=-5 b=1 y=-3x+1 so then now do you see why it's 1 not 10 or 11?

OpenStudy (anonymous):

okay that makes alot of sense thank you